# A Bayesian approach to functional regression: theory and computation

### v2

@@ 4.1 Functional linear regression
Figure 2 shows the results for an underlying $L^{2}$-model, which would be our direct competitor and a more representative test for our Bayesian model. In this case the outcome is satisfactory, as for the most part our models are on a par with the rest, even defeating other methods that were designed with the $L^{2}$-model in mind. Especially interesting is the comparison with the standard $L^{2}$ functional regression (flin in the graphics), which we outperform most of the time.
@@
Figure 3 depicts the results for the real data sets, where we can see that the performance of our one-stage methods is about the same as that of the comparison methods. However, our variable selection methods are somewhat worse that the reference methods, although not by a wide margin. We have to bear in mind that real data is more complex and noisy than simulated data, and it is possible that after a suitable pre-preprocessing we would have obtained better results. Nonetheless, our goal was to perform a general comparison with a uniform methodology, without focusing too much on the specifics of any particular data set.
On another note, we see that some of our Bayesian models have a higher standard deviation, partly because there is an intrinsic randomness in the sampling mechanism, and it can be the cause of the occasional worse performance. In relation to this, we observe that the methods that use the trimmed mean as a summary statistic tend to have a worse score, as this statistic is more sensitive to outliers than the median or the mode.
@@ 4.2 Functional logistic regression
Subsequently, Figure 5 shows that in the $L^{2}$ scenario the results are again promising, since our models score consistently on or above the mean of the reference models, and in many instances surpass most of them. In this case we also beat the alternative functional logistic regression method (flog). In addition, in this situation the overall accuracy of all methods is poor (around 60%), so this is indeed a difficult problem in which even small increases in accuracy are relevant.
@@
As for the real data sets, in Figure 6 we see positive results in general, obtaining in some cases accuracies well above the mean of the reference models. In particular, the weighted sum methods tend to have slightly better results than the MAP methods, which is a trend that was also present in the simulated data sets. This was somewhat expected, since the weighted predictions use the full range of information available.
@@ 5 Conclusion
In this work we have introduced a natural and computationally feasible way of integrating Bayesian inference into functional regression models, by means of a RKHS approach that simplifies the usually hard task of setting a prior distribution on a functional space. The proposed RKHS formulation gives a common framework to all finite-dimensional models based on linear combinations of the marginals of the underlying process, establishing a solid theoretical foundation for these popular models while retaining the functional perspective. Our approximation has the advantage of working with simpler functional parameters, thus increasing the interpretability and ease of implementation, as well as allowing direct comparison of models with different dimension. In addition, this approach works especially well in the logistic case, bypassing the difficulties associated with maximum likelihood techniques and providing a tractable alternative to the more studied methods in the literature.
We have also proved a posterior consistency result that ensures the coherence and correctness of the Bayesian methods we developed. These kinds of results have in other contexts more intricate and restrictive conditions to arrive at essentially the same conclusions as we did, but again the introduction of RKHS’s is the key point to greatly simplifying them. It is worth emphasizing that the methodology introduced in Miller (2023), initially intended for mixtures, can be adapted to a completely different scenario to deal with functional data, giving a positive answer to a problem outside the usual scope of application of Doob’s results.
Lastly, we have presented numerical evidence that supports the proposed Bayesian methodology and its predictive performance with simulated and real data sets. Thanks to our RKHS formulation, we can effectively leverage the capabilities of RJMCMC samplers and integrate the unknown number of components $p$ in the Bayesian procedure, which to our knowledge is a novel application in the FDA context. This practical side of our work goes to show that the prediction methods we constructed from the posterior distribution are competitive against several non-cherry-picked frequentist techniques, especially those based on the usual $L^{2}$-models, while still remaining viable implementation-wise. Of course, we are not suggesting that the $L^{2}$ perspective should be abandoned, but merely offering a simple, theoretically-backed alternative that can perform better in many situations.
@@ Acknowledgments
The authors would like to acknowledge the computational resources provided by the Centro de Computación Científica-Universidad Autónoma de Madrid (CCC-UAM). This research was partially supported by grants PID2019-109387GB-I00 and PRE2020-095147 of the Spanish Ministry of Science and Innovation (MCIN), co-financed by the European Social Fund (ESF). J. R. Berrendero and A. Cuevas also acknowledge financial support from grant CEX2019-000904-S funded by MCIN/AEI/10.13039/501100011033.
@@ A.1 Posterior distributions
For the posterior distributions in our Bayesian formulation, we only compute a function proportional to their log-density, since that is enough for a MCMC algorithm to work. Consider the parameter vector $(p,\theta_{p})$, where $\theta_{p}=(b_{p},\tau_{p},\alpha_{0},\sigma^{2})$ with $b_{p}=(\beta_{1},\dots,\beta_{p})$ and $\tau_{p}=(t_{1},\dots,t_{p})$. Recall that we have a labeled data set of independent observations from $(X,Y)$, denoted by $\mathcal{D}_{n}=\{(X_{i},Y_{i}):i=1,\dots,n\}$. A standard algebraic manipulation in the posterior expression using Bayes’ formula yields the following results.
Proposition 6.
@@
$\log\pi(p,\theta_{p}\mid\mathcal{D}_{n})\propto{}$ $-\frac{1}{2}\left(\frac{\|b_{p}\|^{2}}{\eta^{2}}-\frac{\|\bm{Y}-\alpha_{0}\bm{1}+\mathcal{X}_{\tau_{p}}b_{p}\|^{2}}{\sigma^{2}}\right)-(n+2)\log\sigma$
@@
where $\bm{Y}=(Y_{1},\dots,Y_{n})^{T}$, $\bm{1}$ is an $n$-dimensional vector of ones, and $\mathcal{X}_{\tau_{p}}$ is the data matrix $(X_{i}(t_{j}))_{i,j}$, for $i=1,\dots,n$ and $j=1,\dots,p$.
Proposition 7.
@@
$\log\pi(p,\theta_{p}\mid\mathcal{D}_{n})\propto{}$ $\sum_{i=1}^{n}\left[Y_{i}\left(\alpha_{0}+\left\langle X_{i},\alpha\right\rangle_{K}\right)-\log\left(1+\exp\left\{\alpha_{0}+\left\langle X_{i},\alpha\right\rangle_{K}\right\}\right)\right]$
@@
$-\log\left(100+\alpha_{0}^{2}\right)+\log\pi(p),$
where $\left\langle X_{i},\alpha\right\rangle_{K}=\sum_{j=1}^{p}\beta_{j}X_{i}(t_{j})$.
@@ A.2 Label switching
A well-known issue that arises when using MCMC methods in mixture-like models such as the one proposed in this work is label switching, which in short refers to the non-identifiability of the components of the model caused by their interchangeability. In our case, this happens because the likelihood and the prior are symmetric with respect to the ordering of the parameters $b$ and $\tau$, i.e., $\pi(Y|X,\theta)=\pi(Y|X,\nu(\theta))$ for any permutation $\nu$ that rearranges the indices $j=1,\dots,p$. Thus, since the components are arbitrarily ordered, they may be inadvertently exchanged from one iteration to the next in a MCMC algorithm. This can cause nonsensical answers when summarizing the marginal posterior distributions of the parameters to perform inference, as different labelings might be mixed on each component (Stephens, 2000). This is primarily the reason why we do not directly use summaries of the posterior distribution of the individual parameters in our prediction methods.
@@ A.3 Affine-invariant ensemble samplers
To maintain the affine invariance and the joint distribution of the ensemble, the walkers are advanced one by one following a Metropolis-Hastings acceptance scheme. The authors consider mainly two types of moves.
@@
where $W$ is a normal distribution with mean 00 and the same covariance as the sample covariance of all walkers in $S_{l}$. The acceptance probability in this case is just the Metropolis ratio, namely $\alpha=\min\{1,\pi(\Gamma)/\pi(\Lambda_{l}(t))\}$.
@@ A.4 RJMCMC implementation
For the computational implementation of our Bayesian prediction pipeline we chose the Python library Eryn (Karnesis et al., 2023). This package is a general-purpose suite of MCMC methods which is reliable, easy to use and performs well in a wide class of problems. It is an evolution of the emcee library mentioned in Appendix A.3 that implements affine-invariant ensemble samplers, with a few key improvements:
@@
The main reason for choosing this library is that it implements a reversible jump MCMC sampling scheme, which is an essential element in our proposed models. It allows trans-dimensional sampling for posterior approximation, letting the user select the likelihood, prior and proposal distributions, and providing a great level of control over the details.
@@
This is a mechanism to increase the efficiency with which the sampler explores the parameter space. The basic idea is to consider a set of Markov chains in parallel, each one sampling from a transformed posterior distribution $\pi_{T}(p,\theta_{p}|X,Y)=\pi(Y|X,p,\theta_{p})^{1/T}\pi(p,\theta_{p})$, where $T\geq 1$ is a temperature parameter. In the words of Karnesis et al. (2023), “intermediate temperatures ‘smooth out’ the posterior by reducing the contrast between areas of high and low likelihood”. In practice, these chains periodically exchange information, with the swaps controlled by an acceptance probability that maintains detailed balance, and ultimately we are only interested on the cold chain ($T=1$).
@@
Another advantage of this ensemble approach, apart from the property of affine-invariance, is that it only requires the specification of a few hyperparameters irrespective of the underlying dimension. This contrasts to, say, the $\mathcal{O}(N^{2})$ degrees of freedom corresponding to the covariance matrix of an $N$-dimensional jump distribution in Metropolis-Hastings. We already covered the prior, likelihood and posterior distributions used in Section 2 and in Appendix A.1. We give below an overview of other integral parts of our RJMCMC method with some implementation details.
@@
We need to specify the initial values for the parameters of all the chains in our sampler. In general, we set these values by sampling from the prior distribution of the corresponding parameters. However, in the linear case the prior on $(\alpha_{0},\sigma^{2})$ is improper, so we use ad-hoc weakly informative distributions for the initial values instead. For $\alpha_{0}$ we consider a normal distribution with mean 00 and standard deviation $10|\bar{Y}_{\text{scaled}}|$, where $Y_{\text{scaled}}$ is the version of the original data scaled to have standard deviation unity. For $\sigma^{2}$ we use an inverse-gamma distribution with shape parameter $a=2$ and scale parameter $b=\hat{\sigma}^{2}_{Y}/\operatorname{Var}(Y)$, where $\hat{\sigma}^{2}_{Y}$ is a rough estimate of an acceptable error in the scale of $Y$. This estimation is done is practice as two order of magnitudes less than $|\bar{Y}|$.
@@
For the parameters $\alpha_{0}$ and $\sigma^{2}$, which are common to all sub-models, we use the stretch move explained in Appendix A.3.
@@
For the between-model moves (those which change the dimension), we set an equal probability of births and deaths, except on the end points of the range of $p$. Specifically, $b_{p,p+1}=d_{p,p-1}=0.5$ if $101<p<$ 10, and $b_{1,2}=d_{10,9}=1$. Note that these are only the probabilities to propose the corresponding move, which will be accepted or rejected according to the acceptance formula in Section 2.3. When the birth of a new component is proposed, we use the prior distribution of $b$ and $\tau$ to generate the new values, and when a death is proposed, the method selects an existing component at random as a candidate for deletion.

### v3

@@ 4.1 Functional linear regression
Figure 2 shows the results for an underlying $L^{2}$-model, which would be our direct competitor and a more representative test for our Bayesian model. In this case the outcome is satisfactory, as for the most part our models are on a par with the rest, even surpassing other methods that were designed with the $L^{2}$-model in mind. Especially interesting is the comparison with the standard $L^{2}$ functional regression (flin in the graphics), which we outperform most of the time.
@@
Figure 3 depicts the results for the real data sets, where we can see that the performance of our one-stage methods is about the same as that of the comparison methods. However, our variable selection methods are somewhat worse that the reference methods, although not by a wide margin. We have to bear in mind that real data is more complex and noisy than simulated data, and it is possible that after a suitable pre-preprocessing we would have obtained better results. Nonetheless, our goal was to perform a general comparison with a uniform methodology, without focusing too much on the specifics of any particular data set. On another note, we see that some of our Bayesian models have a higher standard deviation, partly because there is an intrinsic randomness in the sampling mechanism, and it can be the cause of the occasional worse performance. In relation to this, we observe that the methods that use the trimmed mean as a summary statistic tend to have a worse score, as this statistic is more sensitive to outliers than the median or the mode, even with the 10% trimming.
@@ 4.2 Functional logistic regression
Subsequently, Figure 5 shows that in the $L^{2}$ scenario the results are again promising, since our models score consistently on or above the mean of the reference models, and in many instances exceed most of them. In this case we also beat the alternative functional logistic regression method (flog). In addition, in this situation the overall accuracy of all methods is poor (around 60%), so this is indeed a difficult problem in which even small increases in accuracy are relevant.
@@
As for the real data sets, in Figure 6 we see positive results in general, obtaining in some cases accuracies well above the mean of the reference models. In particular, the weighted sum methods tend to have slightly better results than the MAP methods, which is a trend that was also present in the simulated data sets and in the linear regression experiments. This was somewhat expected, since the weighted predictions use the full range of information available for prediction.
@@ 5 Conclusion
We have introduced a natural and computationally feasible way of integrating Bayesian inference into functional regression models, by means of a RKHS approach that simplifies the inherently hard task of setting a prior distribution on an infinite-dimensional space. The proposed RKHS formulation gives a common framework to all finite-dimensional models based on linear combinations of the marginals of the underlying process, establishing a solid theoretical foundation for these popular points-of-impact models while retaining a functional viewpoint. Our approximation relies on simpler functional parameters, thus enhancing interpretability and ease of implementation, and also enables meaningful model comparisons across different dimensions and facilitates the analysis of the effects of local characteristics of the process.
We have also proved posterior consistency results that ensure the coherence and correctness of the Bayesian methods we developed. These kinds of results have in other contexts more intricate and restrictive conditions to arrive at essentially the same conclusions as we did, but again the introduction of RKHS’s is the key point to greatly simplifying them. On the one hand, we have successfully adapted the methodology recently introduced in Miller (2023)—originally intended for mixture models—to the fundamentally different scenario of functional regression, obtaining “Lebesgue”-almost sure consistency for both the model parameters and the unknown number of components. On the other hand, we have derived alternative consistency results and optimal contraction rates through a more sophisticated approach based on Schwartz’s theorem, which is the standard tool in nonparametric consistency problems. In addition, it is worth mentioning that the theorems we have proved are applicable to a wide range of prior distributions, including many that are rather simple and hence easier to work with.
Lastly, we have presented numerical evidence that supports the proposed Bayesian methodology and its predictive performance with simulated and real data sets. Thanks to our RKHS formulation, we can effectively leverage the capabilities of RJMCMC samplers and integrate the unknown number of components into the Bayesian procedure, which to our knowledge is a novel application in FDA. Moreover, a key finding in our empirical studies is that a relatively low number of components is sufficient to obtain good results. This practical side of our work showcases that the Bayesian prediction methods we constructed are competitive against several non-cherry-picked frequentist techniques, especially those based on the usual $L^{2}$-models, while still remaining viable implementation-wise. Of course, our proposed model is not without limitations, and we are not suggesting that the $L^{2}$ perspective should be abandoned, but merely offering a simple, theoretically-backed alternative that can perform better in many situations.
@@ Acknowledgments
The authors would like to acknowledge the valuable computational resources provided by the Centro de Computación Científica-Universidad Autónoma de Madrid (CCC-UAM). This research was partially supported by grants PID2023-148081NB-I00 and PRE2020-095147 of the Spanish Ministry of Science and Innovation (MCIN), co-financed by the European Social Fund (ESF). J. R. Berrendero and A. Cuevas also acknowledge financial support from grant CEX2023-001347-S funded by MICIU/AEI/10.13039/501100011033.
@@ A.1 Posterior distributions
For the posterior distributions in our Bayesian formulation, we only compute a function proportional to their log-density, since that is enough for a MCMC algorithm to work. Consider the parameter vector $(p,\theta_{p})$, where $\theta_{p}=(b_{p},\tau_{p},\alpha,\sigma^{2})$ with $b_{p}=(\beta_{1},\dots,\beta_{p})$ and $\tau_{p}=(t_{1},\dots,t_{p})$. Recall that we have a labeled data set of independent observations from $(X,Y)$, denoted by $\mathcal{D}_{n}=\{(X_{i},Y_{i}):i=1,\dots,n\}$. A standard algebraic manipulation in the posterior expression using Bayes’ formula yields the following results.
Proposition A.1.
@@
$\log\Pi_{n}(p,\theta_{p}\mid\mathcal{D}_{n})\propto{}$ $-\frac{1}{2}\left(\frac{\|b_{p}\|^{2}}{\eta^{2}}-\frac{\|\bm{Y}_{n}-\alpha\bm{1}_{n}+\mathcal{X}_{\tau_{p}}b_{p}\|^{2}}{\sigma^{2}}\right)-(n+2)\log\sigma$
@@
where $\bm{Y}_{n}=(Y_{1},\dots,Y_{n})^{T}$, $\bm{1}_{n}$ is an $n$-dimensional vector of ones, and $\mathcal{X}_{\tau_{p}}$ is the data matrix $(X_{i}(t_{j}))_{i,j}$, for $i=1,\dots,n$ and $j=1,\dots,p$.
Proposition A.2.
@@
$\log\Pi_{n}(p,\theta_{p}\mid\mathcal{D}_{n})\propto{}$ $\sum_{i=1}^{n}\left[Y_{i}\left(\alpha+\left\langle X_{i},\beta\right\rangle_{K}\right)-\log\left(1+\exp\left\{\alpha+\left\langle X_{i},\beta\right\rangle_{K}\right\}\right)\right]$
@@
$-\log\left(100+\alpha^{2}\right)+\log\Pi(p),$
where $\beta(\cdot)=\sum_{j=1}^{p}\beta_{j}K(t_{j},\cdot)$ and $\left\langle X_{i},\beta\right\rangle_{K}=\sum_{j=1}^{p}\beta_{j}X_{i}(t_{j})$.
@@ A.2 Label switching
A well-known issue that arises when using MCMC methods in mixture-like models such as the one proposed in this work is label switching, which in short refers to the non-identifiability of the components of the model caused by their interchangeability. In our case, this happens because the likelihood and the prior are symmetric with respect to the ordering of the parameters $b$ and $\tau$, i.e., $\mathcal{L}(Y|X,\theta)=\mathcal{L}(Y|X,\theta[\nu])$ for any permutation $\nu$ that rearranges the indices $j=1,\dots,p$. Thus, since the components are arbitrarily ordered, they may be inadvertently exchanged from one iteration to the next in a MCMC algorithm. This can cause nonsensical answers when summarizing the marginal posterior distributions of the parameters to perform inference, as different labelings might be mixed on each component (Stephens, 2000). This is primarily the reason why we do not directly use summaries of the posterior distribution of the individual parameters in our prediction methods.
@@ A.3 Affine-invariant ensemble samplers
To maintain the affine invariance and the joint distribution of the ensemble, the walkers are advanced one by one following a Metropolis-Hastings acceptance scheme. The authors consider mainly two types of moves:
@@
where $W$ is a normal distribution with mean 0 and the same covariance as the sample covariance of all walkers in $S_{l}$. The acceptance probability in this case is just the Metropolis ratio, namely $\alpha=\min\{1,\pi(\Gamma)/\pi(\Lambda_{l}(t))\}$.
@@ A.4 RJMCMC implementation
For the computational implementation of our Bayesian prediction pipeline we chose the Python library Eryn (Karnesis et al., 2023). This package is a general-purpose suite of MCMC methods which is reliable, easy to use and performs well in a wide class of problems. It is an enhancement of the emcee library mentioned in Appendix A.3 that implements affine-invariant ensemble samplers, with a few key improvements:
@@
The main reason for choosing this library is that it implements a reversible jump MCMC sampling scheme, which is an essential element in our proposed methodology. It allows trans-dimensional sampling for posterior approximation, letting the user select the likelihood, prior and proposal distributions, and providing a great level of control over the details.
@@
This is a mechanism to increase the efficiency with which the sampler explores the parameter space. The basic idea is to consider a set of Markov chains in parallel, each one sampling from a transformed posterior distribution given by $\pi_{T}(p,\theta_{p}|X,Y)=\pi(Y|X,p,\theta_{p})^{1/T}\pi(p,\theta_{p})$, where $T\geq 1$ is the temperature. In the words of Karnesis et al. (2023), “intermediate temperatures ‘smooth out’ the posterior by reducing the contrast between areas of high and low likelihood”. In practice, these chains periodically exchange information, with the swaps controlled by an acceptance probability that maintains detailed balance, and ultimately we are only interested on the cold chain ($T=1$).
@@
Another advantage of this ensemble approach, apart from the property of affine-invariance, is that it only requires the specification of a few hyperparameters irrespective of the underlying dimension. This contrasts to, say, the $O(N^{2})$ degrees of freedom corresponding to the covariance matrix of an $N$-dimensional jump distribution in Metropolis-Hastings. We already covered the prior, likelihood and posterior distributions used in Section 2 and in Appendix A.1. We give below an overview of other integral parts of our RJMCMC method with some implementation details.
@@
We need to specify the initial values for the parameters of all the chains in our sampler. In general, we set these values by sampling from the prior distribution of the corresponding parameters. However, in the linear case the prior on $(\alpha,\sigma^{2})$ is improper, so we use ad-hoc weakly informative distributions for the initial values instead. For $\alpha$ we consider a normal distribution with mean 0 and standard deviation $10|\bar{Y}_{\text{scaled}}|$, where $Y_{\text{scaled}}$ is the version of the original data scaled to have standard deviation unity. For $\sigma^{2}$ we use an inverse-gamma distribution with shape parameter $a=2$ and scale parameter $b=\hat{\sigma}^{2}_{Y}/\operatorname{Var}(Y)$, where $\hat{\sigma}^{2}_{Y}$ is a rough estimate of an acceptable error in the scale of $Y$. This estimation is done is practice as two order of magnitudes less than $|\bar{Y}|$.
@@
For the parameters $\alpha$ and $\sigma^{2}$, which are common to all sub-models, we use the stretch move explained in Appendix A.3.
@@
For the between-model moves (those which change the dimension), we set an equal probability of births and deaths, except on the end points of the range of $p$. Specifically, $b_{p,p+1}=d_{p,p-1}=0.5$ if $1<p<10$, and $b_{1,2}=d_{10,9}=1$. Note that these are only the probabilities to propose the corresponding move, which will be accepted or rejected according to the acceptance formula in Section 2.3. When the birth of a new component is proposed, we use the prior distribution of $b$ and $\tau$ to generate the new values, and when a death is proposed, the method selects an existing component at random as a candidate for deletion.